OpenStudy (hayleyc24):

2x^4 + 2x^2 +32 can you break this down any farther? Help will be rewarded

OpenStudy (cwrw238):

well first take 2 out = 2(x^4 + x^2 + 16)

OpenStudy (hayleyc24):

why would you take the 2 out its attached to the x squared

OpenStudy (cwrw238):

2 is the GCF of 2, 2 and 32

OpenStudy (cwrw238):

but the expression in the parentheses won't factor

OpenStudy (hayleyc24):

you've lost me and I just realized I forgot to include this f(g(x)) = and then the problem

OpenStudy (hayleyc24):

I've already solved for most of the problem would it help if I gave you the original problem?

OpenStudy (cwrw238):

lets have the whole problem then

OpenStudy (cwrw238):

yes

OpenStudy (hayleyc24):

okay give me a sec

OpenStudy (hayleyc24):

Let f(x) = x2 + x + 2 and g(x) = 2x2 + 5. Find f(g(x)). Show each step of your work.

OpenStudy (cwrw238):

oh ok what you do is replace very occurrence of x in f(x) by g(x)

OpenStudy (hayleyc24):

Plug in the g(x) equation where all the x's are in f(x). (2x^2 + 5)^2 + (2x^2 + 5) + 2 2x^4 + 5^2 + 2x^2 + 5 + 2 2x^4 + 25 +2x^2 + 7 2x^4 + 2x^2 + 32

OpenStudy (cwrw238):

f(x) = x2 + x + 2 f((g(x)) = (2x^2 + 5)^2 + 2x^2 + 5 + 2 = 4x^4 + 20x^2 + 25 + 2x^2 + 7 = 4x^4 + 22x^2 + 32

OpenStudy (cwrw238):

you didnt expand (2x^2 + 5)^2 correctly

OpenStudy (hayleyc24):

Thank you, so the 22 and 4 cannot be combined becuase of the different exponents right?

OpenStudy (cwrw238):

right

OpenStudy (hayleyc24):

Thank you thank you thank you thank you!!!!!!!!

OpenStudy (cwrw238):

yw lol